OpenStudy (anonymous):

Are these monomials? If so, what degree are they? a) -3/7mn^2p^5 b) 5a^2b/8c^3d^2

OpenStudy (kittiwitti1):

\[-\frac{3}{7}mn^{2}p^{5}\text{ and }\frac{5a^{2}b}{8c^{3}d^{2}}?\]

OpenStudy (anonymous):

Yes

OpenStudy (usukidoll):

A monomial is a polynomial with one term example \[3y^2 \]

OpenStudy (usukidoll):

a binomial is a polynomial with two terms example \[3y^2+x \]

OpenStudy (usukidoll):

a trinomial is a polynomial with three terms Example \[3y^2+x+2\]

OpenStudy (anonymous):

So both are monomials?

OpenStudy (usukidoll):

yeah... because both of them are one term. Now we need to find the degree.

OpenStudy (anonymous):

The first one has a degree of 8?

OpenStudy (usukidoll):

ummmmmm whoa. the degree is the variable with the highest number

OpenStudy (anonymous):

I thought you add the exponents

OpenStudy (anonymous):

so its 5?

OpenStudy (usukidoll):

yeah you add the exponents...sorry about that. it does become 8 for the first one because it's the sum of the variables. so that first one was 1+2+5 = 8. sorry about that ^^

OpenStudy (anonymous):

Its okay. For the second one i got -2 but can the degree be negative?

OpenStudy (usukidoll):

2+1-3-2 = 3-3-2 = 0-2 = -2 I got that as well the degree can't be negative though for degree... it needs to be a positive number.

OpenStudy (usukidoll):

I'm probably repeating words because my glasses are dirty and I need to clean them :/

OpenStudy (anonymous):

So what do we do about the negative degree?

OpenStudy (pawanyadav):

Because (a) has degree 2 with highest number b=1,c=-3 and d=-2

OpenStudy (usukidoll):

if I re-write the second monomial \[\frac{5}{8}a^{2}bc^{-3}d^{-2}\] there's negative exponents so it's not a polynomial at all... not a monomial on this one.

OpenStudy (pawanyadav):

Yes ,second one is not a monomial

OpenStudy (usukidoll):

negative degrees :/

OpenStudy (usukidoll):

a BIG NO NO because we need a non-zero exponent for polynomials ... and that applies to monomials, binomials, and trinomials as well

OpenStudy (anonymous):

Ok i get it thank u!